OpenStudy (rosepedal1667):

a woman bought some large frames for $15 and some small frames for $6 each. She bought a total of 19 pair and she paid $150 total how many did she buy of each?

OpenStudy (anonymous):

need the total money she has to pay.

OpenStudy (mathmate):

19 pairs mean 19 of each, isn't it?

OpenStudy (mathmate):

Are they eye-glass frames?

OpenStudy (anonymous):

Let the no of large frames bought be X Then the no of smaller frames bought will be (19 - X ) Cost incurred = 15X + 6( 19 - X) = ? 15X + 114 - 6X = ? 9X = ?-114 X = ?? ANSWER X large frames and x small frames

OpenStudy (anonymous):

need to know total amount spent to finish the rest